MICROWAVE AND SUBMILLIMETER WAVE MEASUREMENTS OF HDCO IN THE ν_{4}, ν_{5} AND ν_{6} BANDS

\underline{A. PERRIN}, J.-M. FLAUD, Laboratoire de Photophysique Moléculaire, CNRS, Université Paris Sud, Campus d’Orsay, Bat 350, 91405 Orsay Cedex, France; J.DEMAISON, L.MARGULÈS, Laboratoire PhLAM, CNRS, Université de Lille I, Bat. P5, 59655 Villeneuve d’Ascq Cedex, France; H. MÄDER, S.WÖRMKE, Institut für Physikalische Chemie, Olshausenstrasse 40, D-24098 Kiel, Germany.

The rotational spectrum of HDCO in the 4^{1}, 5^{1} and 6^{1} excited vibrational states has been investigated in Lille and Kiel using a sample enriched in D. In Lille the measurements were performed in the millimeter (160-600 GHz) region. The spectra in Kiel were recorded using Fourier transform microwave spectrometers in the regions of about 8-18 GHz and 18-26 GHz, employing a rectangular waveguide of 12 m length and a circular waveguide of 36 m length, respectively. These results were combined with the 4^{1}, 5^{1} and 6^{1} infrared energy levels ^{a} obtained in a previous analysis of FTS spectra of the ν_{4} band (CHD bend), ν_{5} band (CHD rocking) and ν_{6} band (out of plane bend) recorded in the 10 μm region at Giessen in order to get the best possible Hamiltonian constants. The energy levels calculation of the 4^{1}, 5^{1} and 6^{1} interacting states accounts for the usual A- and B-type Coriolis resonances in the 5^{1} - 6^{1} and 6^{1} - 4^{1} off diagonals blocks. On the other hand, in order to reproduce the observed microwave transitions and infrared energy levels, it was necessary to use a {J_{x}, J_{z}} non orthorhombic term in the 5^{1} and 6^{1} v- diagonal blocks of the Hamiltonian matrix. Therefore, HDCO exhibits an example of vibrational induced rotational axis switching.

aA. Perrin, J.-M.Flaud, M.Smirnov, and M.Lock, J. Mol. Spectrosc. \underline{203}, 175 (2000).